The speed of a boat in still water is 10mph. The boat travels 24 miles upstream and back downstream in a total of 5 hours. What is the speed of the current?
Found 2 solutions by jorel555, ikleyn:
Answer by jorel555(1290)

If the boat travels 24 miles both ways, then:
24/10-r+24/10+r=5
24(10+r)+24(10-r)=500-5r^2
240+24r+240-24r=500-5r^2
20=5r^2
4=r^2
r=2
The speed of the current is 2 mph!!!!!!!!!

Answer by ikleyn(52776)

.
The speed of a boat in still water is 10mph. The boat travels 24 miles upstream and back downstream in a total of 5 hours. What is the speed of the current?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

It is a typical Upstream and Downstream round trips word problem.

You can find similar fully solved similar problems on upstream and downstream round trips with detailed solutions in the lessons 

    - Wind and Current problems 

    - More problems on upstream and downstream round trips 

    - Selected problems from the archive on the boat floating Upstream and Downstream 


Read them attentively and learn how to solve this type of problems once and for all.